OpenStudy (anonymous):

find the value of the discriminant and describe the number and type of roots x^2+5x+6=0

OpenStudy (mertsj):

What does it mean by discriminant?

OpenStudy (anonymous):

what do you mean? im honestly not sure it just gave me problems n answers to choose from since the darn powerpoint wont come up

OpenStudy (mertsj):

That is your problem. Type "discriminant of quadratic equation" into a search bar and find out what it is.

OpenStudy (anonymous):

For a given quadratic \(ax^2+bx+c\), the discriminant is \(b^2-4ac\). If it's positive, then you have two real roots. Negative, two complex roots. Zero, one real root.

OpenStudy (anonymous):

soo.. do it like uhh 5^2-4(1)(6)=0??

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

well what about 3x^2-5x+6=0....the answers given isnt what i came up with so i know ive done something wrong just done know what....i got -97

OpenStudy (mertsj):

The problem said to find the number and type of roots.

OpenStudy (mertsj):

What did you get for b ^2-4ac?

OpenStudy (anonymous):

D=1

OpenStudy (anonymous):

which is was one of the answer choices and also two rational roots

OpenStudy (mertsj):

yes. And 1 is positive If it's positive, then you have two real roots. Negative, two complex roots. Zero, one real root.

OpenStudy (anonymous):

well got them right but the other 3 idk what im doing wrong, 6x^2-x-12=0

OpenStudy (mertsj):

What is b^2-4ac for that problem?

OpenStudy (anonymous):

x^2-4(6)(12) right?

OpenStudy (mertsj):

Looks to me like the 12 is negative.

OpenStudy (anonymous):

yea sorry forgot to say that but i got 288 which aint one of my choices D=289, two irrational roots or D=289 two rational roots

OpenStudy (mertsj):

1^1-4(6)(-12)=1+288=289

OpenStudy (anonymous):

ooo so i have to put 1 i cant just keep x

OpenStudy (mertsj):